OpenStudy (anonymous):

Which equation could be used to solve this problem? The sum of two consecutive odd integers is 100. Find the integers. A.2n + 3 = 100 B.2n + 1 = 100 C.n + 4 = 100 D.2n + 2 = 100

OpenStudy (anonymous):

@Michele_Laino i'm thinking A

terenzreignz (terenzreignz):

How do you convey "two consecutive odd integers"? I mean, if say, n is an odd integer, what's the next one?

OpenStudy (michele_laino):

given a natural number n, then the number 2n is even number and its consecutive odd numbers are: 2n-1 and 2n+1

OpenStudy (anonymous):

ok so then wouldn't it be B.

terenzreignz (terenzreignz):

Process of elimination? lol

terenzreignz (terenzreignz):

Let's be systematic here :D 3,5 9,11 21,23 all pairs of consecutive odd numbers. In all cases, you did *something* to the smaller odd number to get the bigger. Now, if n was the smaller odd integer, what is the next odd integer?

OpenStudy (anonymous):

hmmmm i'm confused??

terenzreignz (terenzreignz):

To get 5, you add 2 to 3 To get 11, you add 2 to 9 To get 23, you add 2 to 21 In all cases, you always did the same thing to the odd number to get the next consecutive odd number. So, if n was an odd number, what do you do to it to get the next odd number? Come on... you can do this :D

OpenStudy (anonymous):

ohhh ok hld on then.

OpenStudy (anonymous):

7,13,25

OpenStudy (anonymous):

9,15,27

OpenStudy (anonymous):

guys??

terenzreignz (terenzreignz):

Oh, you. What I meant was if you have a smaller odd integer to get the next odd integer, you simply ADD 2 to it. Given that, if n is the smaller odd integer, what's the next odd integer?

OpenStudy (anonymous):

yea i did do that

terenzreignz (terenzreignz):

I didn't mean "add 2" to those example numbers I gave you. I meant add 2... to n. Come one... this is basically spoonfeeding now lol ^^ Peace

terenzreignz (terenzreignz):

Come on*

OpenStudy (anonymous):

ohh

OpenStudy (anonymous):

well ive been thinking D the who'll time

OpenStudy (anonymous):

x + (x+2) = 100

terenzreignz (terenzreignz):

*shrug* you were thinking A and B just earlier. hardly "the whole time" haha

OpenStudy (welshfella):

lets look at 2n + 2 n could be an odd number and 100-2 = 98 so ?

terenzreignz (terenzreignz):

n but close enough. n + (n+2) = 100 just simplify

OpenStudy (anonymous):

then it has to be B then right

OpenStudy (welshfella):

no not B

OpenStudy (anonymous):

ohh no sorry A

OpenStudy (welshfella):

no i think you are guessing

terenzreignz (terenzreignz):

*twitches*

OpenStudy (welshfella):

- look at my last post - we were looking at 2n + 2 = 100

terenzreignz (terenzreignz):

choice C is offended about not being considered XD but yeah, it does seem like you're guessing. tsk... unhealthy habit, that, and you were so close to getting the answer too :|

OpenStudy (anonymous):

no i think you suck @welshfella and thank you very much @terenzreignz

terenzreignz (terenzreignz):

Thank me later when you get the answer! D: let's start with n + (n+2) = 100 Just simplify it!